OpenStudy (anonymous):

Solve the system of equations. Enter your answer as an ordered pair. 3x + 6y = 36 3x - 6y = 0

OpenStudy (chaise):

3x + 6y = 36 3x - 6y = 0 Plus your equations together to elimate the y 6x = 36 x = 6 3x+6y=36 3x3 +6y = 36 9 + 6y = 36 6y = 27 y = 27/6 y = 9/2 (3, 4.5)

OpenStudy (radar):

Change 3 to 6.

OpenStudy (radar):

(6,4.5)

OpenStudy (chaise):

Nooo. Very sorry, the change in x will change your value for y. Let me fix it up

OpenStudy (radar):

Sorry, didn't realize you used different value of x to solve for y.

OpenStudy (chaise):

3x + 6y = 36 3x - 6y = 0 Plus your equations together to elimate the y 6x = 36 x = 6 3x+6y=36 3x6 +6y = 36 18 + 6y = 36 6y = 18 y=3 (3,3) Very sorry, very amatuer mistake. Excuse myself.

OpenStudy (radar):

Now is the "good answeer" lol